OpenStudy (anonymous):

Suppose that a ball is dropped from the upper observation deck of the CN Tower, 450m above the ground. a) What is the velocity of the ball after 5 seconds? B) How fast is the ball when it hits the ground? Thanks :)

OpenStudy (anonymous):

For this problem, I'm converting meters to feet as I am using -32 ft/sec^2 as the gravitational constant, so: 1 Meter = 3.2808399 Feet 450 m = 1476.377955 feet The position function, using the standard gravitational constant at sea level, is: s(t) = -16t^2 + (V0)t + (S0) and the velocity function is: v(t) = -32t + (V0) At t=5, you have for velocity: v(5) = -32(5) + (0) V0 is 0 because the ball is dropped For the ball hitting the ground: s(t) = 0 = -16t^2 + (V0)t + 1476.377955 0 = -16t^2 + (0)t + 1476.377955 0 = -16t^2 + 1476.377955 -> t = sqrt(1476.377955/16) = 9.6059 v(9.6059) = -32(9.6059)

OpenStudy (anonymous):

can velocity be negative ??

OpenStudy (anonymous):

Sure! Velocity is a vector, it has magnitude and direction. Speed has magnitude only and is expressed always as positive.

OpenStudy (anonymous):

Speed = |velocity|

OpenStudy (anonymous):

I did this problem in feet per sec instead of meters per second. Is that ok for you? If not, then you have to re-convert the velocities back to meters/sec. Can you do that?

OpenStudy (anonymous):

yes that is okay :) Thank you

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @joselin12